Gas systems weekly project meeting
Gas systems 2026-10-15
Key Outcomes
The meeting addressed recurring buffer pressure alarms on the RPC system, mass flow controller issues requiring hardware swaps, and motherboard replacement tracking. Key decisions were made on borrowing a high-flow mass flow controller from ALICE to cover ATLAS-RPC needs, and labeling racks with physical identifiers. Gas bottle inventory management and ventilation problems in the lab256 were flagged as ongoing operational concerns.
Decisions Made
· ATLAS RPC PID tuning: KC parameter on ATLAS RPC exhaust MFC changed from 1 to 5 and kept at that value to address slow pressure response.
· ATLAS RPC Mass flow controller swap: High-flow MFC to be taken from ALICE to replace ATLAS RPC's unit which range is too big now; a new one to be ordered as a permanent solution.
· Rack labeling: Physical, permanent labels to be applied to all racks leaving the workshop (→ move on with respect Miguel and Michael presentation on this topic), with numbers cross-referenced in safety valve records.
· R134a distillation Isobutane/SF6 lines for distillation system testing some connections can be prepared ahead of the RPC stop, as those gases are not currently in use.
· Gas bottle procedure: Existing procedure confirmed as workable; Patrick to prepare 2 slides on the process for the next meeting.
Completed
· CMS piping and flushing expected to finish next week; electrical commissioning to follow with Patrick.
· Motherboard numbering system introduced to track replacements and failure reasons.
· CLOUD (Louis-Philippe) Documentation being prepared mapping each mass flow controller to its screen location and physical zone for maintenance reference.
Blockers
· Buffer pressure alarms at 3:30 AM — root cause unclear; candidates are chamber disconnections and CO2 composition (~95%) affecting exhaust regulation.
· Ventilation/temperature in the lab256 remains unresolved; liquid condensation in mass flow controllers and spectrometer is causing repeated downtime.
· MDT pump issue — intermittent failure not yet solved; switching between Pump1 and Pump2 has been a workaround, but underlying fault unresolved.
· CMS cable box blocking ME0 installation position; responsibility for removal disputed between teams (CMS or Gas Team).
Pending Confirmation
· Whether MID will need their MFC back within the year (estimated ~2 years free, but not guaranteed).
· TRT system fate (keep, partial, or discard) — decision rests with universities, not the team.
· CO2 pressure regulator incident at 904 GEM — Abdel to be contacted to clarify whether he touched the regulator; communication protocol with other experiments to be reinforced.
· CMS RPC Econex valve intervention timing — originally planned for December 2026 AUG test, but very likely it was already postponed to 2027. Confirmation needed by Andrea/Maria Cristina. Gianluca's software readiness unconfirmed.
Risks/Escalations
· ATLAS RPC input flow already at minimum (~25 m³/h); further chamber disconnections may require flow reduction below operable threshold.
· ATLAS RPC Two circulation pumps not functioning correctly — one may need to be sent back to the manufacturer for inspection; test bench for pump validation still not set up.
· Gas bottle rental costs difficult to track due to poor software usability; team losing time on administrative overhead.
· 904 Interventions by other groups (nitrogen, CO2) affecting adjacent systems without prior notification — recurring issue with GEM/RPC panels.
Action Items
· Davide/Max: Investigate buffer pressure spike at 3:30 AM — determine if cause is chamber disconnection or other (e.g., CO2 composition effect).
· Davide/Louis-Philippe: Organize MFC swap from ALICE to RPC as soon as possible given ongoing chamber disconnections.
· Patrick: Prepare 2 slides on gas bottle registration procedure for next meeting.
· Team: Compile full bottle inventory and add entry for all purchased bottles to tracking system.
· Gianluca: Resume functional analysis for SciFi/NDT pump stop logic and confirm with experiment whether power-cut behavior is intended.
· David: Follow up on suspected purifier leak once unit is back in use (January).
· Team: Label all racks physically and retrofit existing racks with identifiers; update safety valve records with rack numbers.
· Team: Open/reopen ventilation ticket daily until resolved.
· LS3 planning: Begin ordering material for RPC; update planning table once delivery timeline is known.
Gas systems 2026-10-07
Action Items
· Max/Davide: Organize meeting with Paolo (Atlas muon coordinator) to escalate Nitrogen incident.
· Roberto: Send updated schedule reflecting holiday period (Jan 3–5 holidays, Jan 6–7 working days, Jan 7 = first Thursday).
· Patrick / Team: Assist with gas bottle registration and return. Prepare few slides with procedure
· Team: Compile full gas bottle inventory list for next week; identify bottles to return vs. keep.
Gas systems 2026-10-01
Action Items
· Gianluca: Verified buffer pressure alarm delay unit (5 seconds vs. 5 minutes) in recipe.-> there is an issue in the recipe -> checked again, it's in seconds (default 5 s)
· Patrick/Stefania/Gianluca: Clarify LHCb UT pipe/valve condition item; check functional analysis for manual stop behavior.
· NA62 Straw: Fabio available for up to two days; Roberto to confirm based on bottle availability.
Gas systems 2026-09-24
Action Items
· Gianluca: Investigate PLC logic for minimum valve opening threshold below 10%. CMS-RPC humidifier -> no values clipped in the PLC found
· Stefania (microcirculation): Order mechanical and electrical material; report estimated arrival at next meeting.
· Davide/Max/Gianluca: Check historical pump test record to identify which pump had the prior current absorption problem ATL-MDT. We tested VFD of pump 1 with pump 2 and it seemed ok: see https://eloggas.web.cern.ch/GASSYSOperationMaintenance/6610
Gas systems 2026 09 17
Action Items
· Davide/Max: Continue CO2 pressure-drop tests on Column B (Atlas-RPC) — check if drop rate decreases over repeated tests (saturation indicator). -> stable rate
· Davide: Perform full pump maintenance on all three RPC pumps (gasket material upgrade, needle valve replacement). -> done maintenance p2, non return completely broken. replaced. pump 2 still not ok after. Plan to replace all non-return -> something solved, but pump still not happy
Gas systems 2026-09-03
Action Items
Patrick/Gianluca: Investigate LHCb-UT, confirm correct logic and pipe (or valve or pneumatic pipes?) condition.- Andrea : finalize TID 904 and send link.
- Florian's mixer: Order required materials — planned for next week – who is ordering? Patrick for electrical
piquet:
LHCb pump filters to be changed or something (for LS3)-> TICKET
O2+H2O analyser to be checked -> TICKET
ATLAS-RPC rack 66 and 69 pressure alarm during purif 2 change over -> TICKET
NA62
glitch of CEDAR 6105, is it the keller? -> TICKET to check
ATLAS-TGC putifier Col B leak. loosing 1 bar in 1 day -> TICKET Davide check with Argon or CO2
CMS-DT pressure sensor PT1303 went to 0 for 20 minutes -> TICKET to follow Andrea
ATLAS-RPC purif -> affecting chamber pressure -> R66 check
12/2 ATL-RPC purif 1 some software issue and sms not indicating the right reason for the alarm -> TICKET
IR status, spares, .... --> TICKET
test binary gas analyser... -> TICKET
outcome of binary gas analyser test --> TICKET & buy a 2nd one --> Maria Cristina
DP pressure for pump filter in ALICE and LHCb -> TICKET -> @LPh + @Patrick [GASSYSTEMS-903] DP pressure for pump filter increasing - CERN Central Jira -> Stefania
chiller fluids to uniform between CMS, LHCb and ATLAS -> TICKET Maria Cristina
check if NV in purif is really useful --> TICKET
Calibration bottles O2-H2O status
ALICE
70 report from Anité [GASSYSTEMS-70] TPC gas quality - filters to Benoit - CERN Central Jira (I don't see the report, can you please add it?)
several alarms blocked: [GASSYSTEMS-628] ALICE alarms blocked - CERN Central Jira
TOF gas mixture analysis -> [GASSYSTEMS-639] TOF gas mixture analysis - CERN Central Jira & ELOG
TRD membrane is connected ? -> [GASSYSTEMS-717] TRD membrane status - CERN Central Jira
P&ID drawing check who is doing -> [GASSYSTEMS-736] ALICE P&ID drawing check - CERN Central Jira
MID swap of pumps during MD because sometime the system stop -> TICKET
ATLAS
MDT frequency converter (need to be replaced TICKET) ELOG - ATLAS MDT Test convertisseur 4101 avec la pompe 4201
RPC distribution rack regulation valve (TICKET)
RPC secondary mixer (TICKET), info to be given to ATLAS
RPC purif 3 circuit breakers & meeting -> [GASSYSTEMS-718] RPC purif 3 electrical issue - CERN Central Jira --> meeting when?
ATLAS-TGC flow difference --> [GASSYSTEMS-612] ATLAS-TGC flow difference - CERN Central Jira
CMS
RPC purif test at workshop, PLC needed -> TICKET
R134a insulation - isol'tut to check and see what to do -> ticket IO [GASSYSTEMS-578] R134a insulation - isol'tut to check and see what to do - CERN Central Jira
How much gas flow we need for the RPC purifier? --> [GASSYSTEMS-579] CMS-RPC How much gas flow we need for the RPC purifiers? - CERN Central Jira @MariaCristina notes to be added
RPC bypass pump too much closed --> [GASSYSTEMS-604] RPC bypass pump too much closed - CERN Central Jira
RPC with CO2 like ATLAS and effect on R134a recuperation -> [GASSYSTEMS-606] RPC with CO2 like ATLAS and effect on R134a recuperation - CERN Central Jira
R134a Upgrade (hardware and software) to be tested -> [GASSYSTEMS-594] R134a Upgrade (hardware and software) to be tested - CERN Central Jira
DT vaisala reading difference between SG and US -> [GASSYSTEMS-677] CMS DT vaisala reading difference between SG and US - CERN Central Jira
CMS-FGR -> TICKET should we add a module. Do we need to buy something? follow-up
CMS-FGR -> TICKET Maria Cristina già fatto per bypass
CMS-FGR -> TICKET flowmeter for Huber
CMS-RPC IR not stable -> elog Maria Cristina -> TICKET
CMS-ID flushing. refurbishing of old racks material and manpower need to be estimated -> TICKET @Gianluca ticket number for project
R134a huber temperature can be controlled by PLC -> TICKET
LHCb
TS1-> RICH1 leak search (postponed), SciFi some work from others expected
RICH1 what we learn from 1st fill -> [GASSYSTEMS-608] RICH1 what we learn from 1st fill - CERN Central Jira
summary of the C4F10 bottles --> TICKET or ELOG -> [GASSYSTEMS-660] summary of the C4F10 bottles - CERN Central Jira
alarms blocked: [GASSYSTEMS-631] LHCb alarms blocked - CERN Central Jira
RICH2 filter blocking -> [GASSYSTEMS-731] RICH2 pump filter differential pressure sensor not working - CERN Central Jira
P&ID drawing check who is doing -> [GASSYSTEMS-739] LHCb P&ID drawing check - CERN Central Jira
RICH1 cleaning , when? not before mid-september? -> TICKET
RICH1 remove membrane rack before end of HL-LHC operation and/or change piping-> TICKET
RICH2 injection through recuperated CF4 -> TICKET
fire detection alarms reported by Pieter, more info? but not on us. Cleaning of SG room -> TICKET
CODEX-b interlock for flammable gas detection -> Patrick
RICH1 leak repair -> December 9th tbc -> TICKET
NA62
CEDAR maintenance needed -> [GASSYSTEMS-609] CEDAR maintenance needed - CERN Central Jira
change of pressure sensor used for leak warning? -> [GASSYSTEMS-741] NA62-CEDAR move to keller pressure sensor? - CERN Central Jira
GIF++ TGC gas system and return flow: [GASSYSTEMS-624] GIF++ TGC gas system return flow - CERN Central Jira
[GASSYSTEMS-674] MFCs to be checked - CERN Central Jira :
- Fondoscala del MFC di R134A e sensore umidità del gas RPC [GASSYSTEMS-645] Fondoscala del MFC di R134A e sensore umidità del gas RPC - CERN Central Jira
- Neutrino platform gas analysis: [GASSYSTEMS-622] RE: Analisi di outgassing di diversi materiali - CERN Central Jira
- ALICE T10 MFCs: [GASSYSTEMS-669] ALICE T10 TR: calibration MFC pour C4F10 - CERN Central Jira
- ND280 MFCs to check [GASSYSTEMS-679] ND280 MFCs to check - CERN Central Jira
non-LHC safety valve --> check lista di Fabio
[GASSYSTEMS-172] non-LHC safety valves to be verified - CERN Central Jira
ELOG - safety valves to be verified for non-LHC systems
Additional requests/projects/discussions
887 CEDAR new "mixers" --> [GASSYSTEMS-509] CEDAR for beam lines - CERN Central Jira
AMBER
EDMS leak test?
test certificate x-ray?
ATEX certificate?
who is technical coordinator
resource manager Oleg D?
no money? should we stop? and resume in 2026?
Commissioning when? -> TICKET
when they would like to start? Moritz? -> ask
LEAK DETECTORS LIST -> TICKET
MAINTENANCE INTERVENTIONS WORK ORDER FOR STATISTICS -> TICKET
PROJECT 2025
summary TICKET
examples of topics for the common discussions:
-cms id flushing -> Andrea
-cms r134a operation and spare -> Andrea
-cms cf4 operation and spare -> Andrea
-atl tgc nc5h12 -> Kacper
-Analysis boxes -> LPh
-C4F10 new recuperation installation -> Patrick